# Sell It Before You Make It: Revolutionizing E-Commerce with Personalized AI-Generated Items

### v1

@@ 4.3. Personalized Image Generation
In this section, we propose the PerFusion framework for personalized group-level preference alignment based on stable diffusion (Rombach et al., 2022). Specifically, as shown in Figure 4, we first design the personalized adaptive network to capture the diverse preference across different users, and then derive the group-level preference optimization objective to model the users’ comparative pattern among multiple candidate images.
@@ 4.3.1. Personalized Adaptive Network
Stable diffusion is essentially a U-Net (Ronneberger et al., 2015) with an encoder, a middle block, and a skip-connected decoder. As shown in Figure 4(a), it contains a total of 25 blocks, including one middle block, and 12 blocks for encoder and decoder, respectively. Each block consists of convolution layers and other necessary neural operators like pooling and batch normalization. The symbol “×$\times 3$” indicates that the corresponding block repeats three times. The U-Net takes the latent representation $z_{t}$ to be denoised, the timestep $t$, and the textual condition $c$ as inputs, and outputs the noise to be eliminated, i.e., $\epsilon_{\theta}(z_{t},t,c)$.
@@ 4.3.2. Group-Level Preference Optimization
We derive the training objective based on DPO (Wallace et al., 2024; Rafailov et al., 2024), the most popular pairwise preference alignment algorithm. Formally, we can divide the candidate image set $\mathcal{X}$ into a positive set $\mathcal{P}$ and a negative set $\mathcal{N}$ according to user binary feedback. Then, we can first substitute the pairwise Bradley-Terry model (Bradley and Terry, 1952) in DPO (Rafailov et al., 2024) with the Plackett-Luce model (Luce, 1959; Plackett, 1975) to consider group-level preference distributions with an independent assumption among positive samples:
@@
[…]
@@
where $r(x,c)$ is the reward model that can be derived according to previous works (Wallace et al., 2024; Rafailov et al., 2024) with a fixed reference model $p_{ref}(\cdot)$:
@@
However, it is intractable to calculate the likelihood of an image, i.e., $p_{\theta}(x|c)$. Therefore, following previous work (Wallace et al., 2024), we approximate the denoising trajectory $p_{\theta}(x|c)$ with the diffusion process $q(x|c)$ and derive the final objective:
@@
5. Experiment
We conduct experiments to answer research questions as follows:
  • RQ1
    How does PerFusionRM perform compared with existing baselines for the user preference estimation task?
  • RQ2
    How does PerFusion perform compared with existing baselines for the personalized image generation task?
  • RQ3
    How do different components and configurations affect the performance of PerFusionRM and PerFusion?
@@ 5.1.1. Datasets
We adopt the public PickaPic dataset (Kirstain et al., 2023) and Alibaba’s Industrial dataset for experiments. The dataset statistics are given in Appendix B.1 due to page limitations.
@@ 5.1.2. Baselines
For user preference estimation task, we adopt Aesthetics (Schuhmann, 2022), CLIP-H (Radford et al., 2021), ImageReward (Xu et al., 2024a), HPSv1 (Wu et al., 2023b), HPSv2 (Wu et al., 2023a) and PickScore (Kirstain et al., 2023) as the baselines. For personalized image generation task, we adopt Stable Diffusion v1.5 (Rombach et al., 2022), SFT (Ho et al., 2020), DPO (Wallace et al., 2024), SPO (Liang et al., 2024), PMG (Shen et al., 2024) as baselines. We give detailed information about baselines in Appendix B.2 due to page limitations.
@@ 5.1.3. Evaluation Metrics
For user preference estimation task, we formulate it as a ranking task to rank positive images with higher scores. Hence, we adopt the mean averaged precision (MAP) and group-wise area under the curve (GAUC) as the metrics. For the personalized image generation task, following previous works (Wallace et al., 2024; Ho et al., 2020; Liang et al., 2024; Xu et al., 2024d) we use Aesthetic (Schuhmann, 2022), CLIP Score (Radford et al., 2021), and PickScore (Kirstain et al., 2023) to evaluate the general quality of generated images. Moreover, we adopt our proposed PerFusionRM as the reward model to estimate whether the generated images can satisfy the users’ personalized preferences. We give detailed information for evaluation metrics in Appendix B.3 due to page limitations.
@@ 5.1.4. Implementation Details
5.2. User Preference Estimation (RQ1)
@@
    • PickaPic is a public dataset, where the candidate images are generated by various open-sourced models like stable diffusion variants (Rombach et al., 2022). The images for text prompts might be imperfect and force users to concentrate more on the text-image consistency and general aesthetic quality. Hence, personalized preference factors are weakened on PickaPic dataset.
@@
5.3. Personalized Image Generation (RQ2)
@@ 6. Industrial Deployment
In this section, we provide a detailed description and discussion about the real-world deployment of personalized image generation for AI-generated items (AIGI) in Alibaba, a world-wide e-commercial platform. Specifically, in our application, the generation mainly focuses on the fashion topics, which can be easily supported for manufacturing solely based on the product design figures. Next, we would like to discuss the industrial deployment from the following four aspects: human expert evaluation, online performance, case study, and broader impact.
@@ 6.1. Human Expert Evaluation
In this section, we aim to evaluate the quality of AI-generated items, and assess how they compare to real human-designed and photographed images. We involve a panel of 8 experts, comprising 6 professional fashion designers and 2 fashion retail specialists, to conduct a blind evaluation of the images, unaware of whether the given image is AI-generated or human-designed. We assess the images from six aspects, including material quality, model realism, craft design, pattern appropriateness, design aesthetics, and color coordination. For each aspect, experts rate the images using a three-tier scoring system: 0 for poor, 1 for average, and 2 for good.
The evaluation results are shown in Table 3. We can observe that AI-generated item images are generally close to human-designed real ones, especially in design aspects such as material quality, pattern appropriateness, and color coordination. While there is still room for improvement in craft design, our practical validation demonstrates that current AI-generated items (AIGI) can already support online e-commercial retail without the need for real-world manufacturing or physical prototypes beforehand.
6.2. Online Performance
Table 4. The relative improvement of click-through rates (CTR) and conversion rate of AI-generated items compared with their non-AI counterparts in the same store.
CTR for Search CTR for recommendation Conversion Rate
Rel.Imrpv. +17.81% +13.35% +17.27%
@@ 6.2. Online Performance
We report the relative improvements of click-through rate (CTR) for both search and recommendation scenarios, as well as the conversion rate, for the AI-generated fashion items compared with human-designed ones. The relative improvement is derived by comparing AI and non-AI items within the same e-commercial store and averaging the results at the store level. For fair comparison, AI-generated items and human-designed items are treated equally by the platform in terms of search and recommendation traffic. There is no preferential treatment towards either product type.
The results are reported in Table 4. We observe that the click-through rate (CTR) for AI-generated items has increased by an average of 15.58% in both the search and recommendation scenarios, while the conversion rate has improved by 17.27%. This highlights the advantages of AI-generated items (AIGI) in attracting user clicks and driving purchase conversions.
We attribute these improvements to two potential factors. First, the AI-generated digital models enhance the aesthetic appeal of the product images, which many smaller fashion shops may struggle to achieve. Second, the rapid iteration capabilities of our AI-driven design process significantly reduce the time needed to launch new products from month-level cycles to day-level cycles. This acceleration eliminates the need for human design, prototype production, and model photography, allowing the items to stay much more aligned with fast-changing fashion trends and consumer preferences, thereby resulting in better engagement and conversion rates.
@@ 6.3. Case Study
We conduct a case study to demonstrate the personalized preferences across different individual users (i.e., the merchants). The construction of the text prompt in our application is simply concatenating different descriptive phrases with commas. To assist users in formulating precise inputs, our system also offers suggestions for professional descriptions. For example, as shown in Figure 7, the designers from two different e-commercial stores submit the same text prompt for fashion product generation, i.e., “dress,v-neck,slim-fit,long sleeves,long-hair model”. Note that during the training of our model, the descriptive phrases can be shuffled for data augmentation, and the prompt is thereby order-insensitive to phrases. From Figure 7, we draw the following observations:
@@
  • More importantly, even when the text prompt is the same for designers from two different e-commercial stores, the resulting images are tailored to reflect the distinct preferences and records of individual users, with a significant difference in general style and design details. We give our analysis as follows:
@@
In summary, the case study demonstrates that our deployed model can not only satisfy the general aesthetic preference and text-image consistency, but also capture the naunced personalized preferences across different individuals.
6.4. Broader Impact
The deployed system in this paper does not only address the key scientific challenge of modeling users’ group-level personalized preferences for text-to-image generation, but also presents a revolutionary potential paradigm for e-commerce under the “sell it before you make it” business mode with AI-generated items (AIGI).
Firstly, our system empowers merchants to respond to the emerging trends and ever-evolving customer preferences faster than ever before, enabling a more dynamic and consumer-responsive e-commerce landscape. By enabling the creation of personalized product designs on demand, we can streamline the preliminary design phases, cut down costs, and speed up the time to market from month-level to day-level.
Secondly, AIGI substantially reduces the need for physical prototypes, minimizing the waste and inventory risk for e-commerce. Only when the number of orders for one AI-generated item exceeds a threshold do the merchants start to produce it. Otherwise, the merchant can choose to offer refunds instead of going into production. In fact, with the fast speed and large volume of new items released via AIGI, merchants can always hit the top-selling products out of the AI-generated items and cover the costs of other refunding ones, thereby achieving nearly zero inventory risk.
@@ 7. Conclusion
This paper presents an innovative system deployed at Alibaba that leverages personalized text-to-image generation for AI-generated items (AIGI), enabling the “sell it before you make it” business mode. We propose a novel PerFusion framework to model the users’ group-level personalized preferences under such scenarios. We conduct both offline and online experiments to validate the effectiveness of our algorithm. The AI-generated items have achieved over 13% improvement in both click-through rate and conversion rate over human-designed ones, highlighting the revolutionary potential of our developed AIGI for e-commercial platforms.
@@ A.1. Ideal Preference Distribution
The classical pairwise preference optimization algorithms such as DPO (Rafailov et al., 2024; Wallace et al., 2024) are derived based on the Bradley-Terry model. In order to consider the preference optmization towards multiple candidates, we have to replace the Bradley-Terry model (Bradley and Terry, 1952) with the Plackett-Luce model (Luce, 1959; Plackett, 1975). Specifically, the Plackett-Luce model probability for a ranking is given by:
@@
[…]
@@
where $\text{Perm}(\cdot)$ is full permutations of a given set, and, according to previous works (Wallace et al., 2024; Rafailov et al., 2024), the reward $r(x,c)$ can be derived as:
@@ A.2. Loss Function
According to previous works on preference optimization (Wallace et al., 2024; Rafailov et al., 2024; Bai et al., 2022), the training of the target model $p_{\theta}(\cdot)$ is to maximize the log likelihood of the ideal ranking for a given set of candidates. We can first derive the log likelihood of the ideal ranking based on the Plackett-Luce model:
@@
Note that it is intractable to calculate the likelihood of an image, i.e., $p_{\theta}(x|c)$. Therefore, following previous work (Wallace et al., 2024), we approximate the denoising trajectory $p_{\theta}(x|c)$ with the diffusion process $q(x|c)$ and derive the final objective:

### v2

@@ 4.3. Personalized Image Generation
We propose the PerFusion framework for personalized group-level preference alignment based on stable diffusion (rombach2022high). As shown in Figure 4, we first design the personalized adaptive network to capture the diverse preference across different users, and then derive the group-level preference optimization objective to model the users’ comparative pattern among multiple candidate images.
@@ 4.3.1. Personalized Adaptive Network
Stable diffusion is essentially a U-Net (ronneberger2015u) with an encoder, a middle block, and a skip-connected decoder. As shown in Figure 4(a), it contains a total of 25 blocks, including one middle block, and 12 blocks for encoder and decoder, respectively. Each block consists of convolution layers and other necessary neural operators like pooling and batch normalization. The symbol $\times 3$” indicates that the corresponding block repeats three times. The U-Net takes the latent representation $z_{t}$ to be denoised, the timestep $t$, and the textual condition $c$ as inputs, and outputs the noise to be eliminated, i.e., $\epsilon_{\theta}(z_{t},t,c)$.
@@ 4.3.2. Group-Level Preference Optimization
We derive the training objective based on DPO (wallace2024diffusion; rafailov2024direct), the most popular pairwise preference alignment algorithm. Formally, we can divide the candidate image set $\mathcal{X}$ into a positive set $\mathcal{P}$ and a negative set $\mathcal{N}$ according to user binary feedback. Then, we can first substitute the pairwise Bradley-Terry model (bradley1952rank) in DPO (rafailov2024direct) with the Plackett-Luce model (luce1959individual; plackett1975analysis) to consider group-level preference distributions with an independent assumption among positive samples:
@@
$=$ $\sum\nolimits_{x_{p}\in\mathcal{P}}p(x_{p}\succ x_{n},\forall x_{n}\in\mathcal{N})\;\;\;\text{(by independent assumption)}$
@@
where $r(x,c)$ is the reward model that can be derived according to previous works (wallace2024diffusion; rafailov2024direct) with a fixed reference model $p_{ref}(\cdot)$:
@@
However, it is intractable to calculate the likelihood of an image, i.e., $p_{\theta}(x|c)$. Therefore, following previous work (wallace2024diffusion), we approximate the denoising trajectory $p_{\theta}(x|c)$ with the diffusion process $q(x|c)$ and derive the final objective:
@@
5. Offline Experiment
@@ 5.1.1. Datasets
We adopt the public PickaPic dataset (kirstain2023pick) and Alibaba’s Industrial dataset for experiments. We show the dataset statistics in Table 1.
Table 1. The dataset statistics.
Dataset #Users #Prompts #Samples Group Size
PickaPic 3,094 33,164 529,219 2
Industrial 54,995 60,616 607,328 5
Crucially, while Pickapic dataset lacks the rich merchant-specific features of our industrial data, it provides a user ID for each preference judgment. This allows our personalized models to learn user-specific tastes, although the more limited feature set explains the comparatively smaller performance gains observed on this dataset.
We filter Pickapic and industrial datasets to ensure each user has at least 10 group-level records. We split datasets into train/valid/test sets with 8:1:1 ratio at group record level, and all evaluations are done on hold-out sets.
@@ 5.1.2. Baselines
For user preference estimation task, we adopt Aesthetics (aesthetic_classifier), CLIP-H (radford2021learning), ImageReward (xu2024imagereward), HPSv1 (wu2023human), HPSv2 (wu2023humanv2) and PickScore (kirstain2023pick) as the baselines. For personalized image generation task, we adopt Stable Diffusion v1.5 (rombach2022high), SFT (ho2020denoising), DPO (wallace2024diffusion), SPO (liang2024step), PMG (shen2024pmg) as baselines. We give detailed information about baselines in Appendix B.1 due to page limitations.
@@ 5.1.3. Evaluation Metrics
For user preference estimation task, we formulate it as a ranking task to rank positive images with higher scores. Hence, we adopt the mean averaged precision (MAP) and group-wise area under the curve (GAUC) as the metrics. For the personalized image generation task, following previous works (wallace2024diffusion; ho2020denoising; liang2024step; xu2024personalized) we use Aesthetic (aesthetic_classifier), CLIP Score (radford2021learning), and PickScore (kirstain2023pick) to evaluate the general quality of generated images. Moreover, we adopt our proposed PerFusionRM as the reward model to estimate whether the generated images can satisfy the users’ personalized preferences. We give detailed information for evaluation metrics in Appendix B.2 due to page limitations.
@@ 5.1.4. Implementation Details
5.2. User Preference Estimation
@@
    • PickaPic is a public dataset, where the candidate images are generated by various open-sourced models like stable diffusion variants (rombach2022high). The images for text prompts might be imperfect and force users to concentrate more on the text-image consistency and general aesthetic quality. Hence, personalized preference factors are weakened on PickaPic dataset.
@@
5.3. Personalized Image Generation
@@ 5.4. Ablation Study
5.5. Computational and Memory Cost
We train our models on 8 NVIDIA GPUs and perform inference on a single GPU. As shown in Table 4, both the training and inference time of PerFusion are higher than the standard diffusion model. This increased overhead is brought by the personalized adaptive network required for processing user-specific features. In our production environment, this trade-off is justified by the significant business impact, including a +17.27% increase in conversion rate and -7.9% decrease in return rate for the apple-to-apple comparison in the same store. This makes the additional cost a worthwhile investment for achieving superior personalization.
As for memory consumption, PerFusion runs with FP16 inference, using  9GB GPU memory (vs. 5GB for SD1.5) due to the personalized adaptive network. In production, we apply techniques like xFormers to further reduce memory usage, making the system scalable for large-scale e-commerce deployment.
Table 4. The computational cost analysis of different models in terms of training and inference time.
Model Train Time/Batch Infer. Time/Sample
Pickapic Industrial Pickapic Industrial
Diffusion 1.45s 1.48s 1.49s 1.49s
DiffusionDPO 3.05s 3.16s 1.48s 1.50s
PerFusion 6.29s 6.95s 3.03s 3.05s
5.6. Backbone Analysis of PerFusionRM
We further conduct experiments by replacing the CLIP backbone of PerFusionRM with other multimodal models:
  • SigLIP (zhai2023sigmoid): A dual-tower model trained with pairwise sigmoid loss over visual and textual encoders.
  • ViLT (kim2021vilt): A single-tower model that jointly captures the relationship of vision and text based on shallow input layers.
Table 5. The backbone analysis of PerFusionRM.
Backbone CLIP SigLIP ViLT
MAP 0.8822 0.8798 0.8776
GAUC 0.7644 0.7601 0.7577
The results are reported in Table 5. CLIP outperforms SigLIP, even though they both use dual-tower architectures. This is mainly because we adopt the PickScore-style CLIP variant, which is first pre-trained on the same dataset to better fit the distribution, leading to more accurate user preference modeling. ViLT, as a single-tower model that fuses vision and text inputs via concatenation, performs relatively worse. A likely reason is that only one personalized plug-in module can be applied to the mutual text-image encoder, which limits its ability to separately model user preferences over vision and text. In contrast, dual-tower models can disentangle these modalities, allowing more fine-grained and flexible personalization.
@@ 6. Industrial Deployment
We provide a detailed description and discussion about the real-world deployment of personalized image generation for AI-generated items (AIGI) in Alibaba, a world-wide e-commerce platform. Specifically, in our application, the generation mainly focuses on the fashion topics, which can be easily supported for manufacturing solely based on the product design figures.
@@ 6.1. Human Expert Evaluation
We aim to evaluate the quality of AI-generated items, comparing to real human-designed and photographed images. Our human evaluation is conducted by a core panel of 8 experts, comprised of 6 professional fashion designers and 2 retail specialists, each with over 10 years of experience in the industry. To ensure quality and scale, annotations from a larger team of outsourced labelers are all reviewed and validated by these experts, guaranteeing a high standard of accuracy and professionalism across the entire evaluation dataset. We assess the images from six aspects, including material quality, model realism, craft design, pattern appropriateness, design aesthetics, and color coordination. For each aspect, experts rate the images using a three-tier scoring system: 0 for poor, 1 for average, and 2 for good.
The evaluation results are shown in Table 6. We can observe that AI-generated item images are generally close to human-designed real ones, especially in design aspects such as material quality, pattern appropriateness, and color coordination. While there is still room for improvement in craft design, our practical validation demonstrates that current AI-generated items (AIGI) can already support online e-commerce retail without the need for real-world manufacturing or physical prototypes beforehand.
6.2. Product Lifecycle & Manufacture Feasibility
A primary concern for our “sell it before you make it” model is its real-world viability, specifically regarding the production lifecycle, manufacturing feasibility, and customer satisfaction. Our deployment at Alibaba has validated this innovative business mode by establishing a seamless and efficient workflow from digital design to physical product. The lifecycle and feasibility are detailed below.
6.2.1. Product Lifecycle
Our AIGI service is a merchant-facing tool, where end customers are generally unaware that a product originated from an AI design. This ensures a natural and uninterrupted shopping experience. The process unfolds in three key stages:
  • AI-Assisted Design: Merchants use our service to generate and select photorealistic designs to be listed in online stores.
  • Demand Aggregation: The AI-generated products appear alongside human-designed items on the platform. Customers browse and place orders as they normally would.
  • Manufacturing: Customer orders are aggregated. Once demand reaches a set threshold, a bulk production order is triggered. The manufactured products are then delivered to end customers.
6.2.2. Manufacturing Feasibility & Customer Satisfaction
A core challenge for an AI-driven, on-demand fashion model is the feasibility of rapid, high-fidelity manufacturing. We validated our system through a large-scale industrial deployment, demonstrating that modern supply chains can support this approach at scale.
Our primary validation comes from a 6-month online A/B test across 17 product categories. We compared the customer return rates of our AI-generated items against human-designed items within the same stores. The AI-generated products exhibited a 7.9% lower return rate, providing strong quantitative evidence of high customer satisfaction. This result indicates that the physical products successfully meet or exceed the expectations set by the AI-generated images. The operational success of this model is underpinned by two key factors of manufacturing facilities:
  • Production Fidelity. The AI-generated images serve as sufficient and direct blueprints for manufacturing, enabling accurate production without physical prototypes.
  • Rapid Turnaround. Production cycles range from 3 to 20 days, with a mean of 7 days, making the on-demand customer experience competitive with traditional in-stock inventory models.
This confirms that our integration of generative AI and responsive manufacturing constitutes a viable and scalable paradigm.
6.3. Post-Launch Performance
Table 7. The relative improvement of click-through rates (CTR), conversion rate and return rate of AI-generated items compared with their non-AI counterparts in the same store.
[…]
Rel.Imrpv. +17.81% +13.35% +17.27% -7.9%
@@ 6.3. Post-Launch Performance
We report relative improvements of click-through rate (CTR) for both search and recommendation scenarios, as well as conversion rate and return rate, for the AI-generated items compared with human-designed ones. The relative improvement is derived by comparing AI and non-AI items within the same e-commerce store and averaging the results at the store level. For fair comparison, AI-generated items and human-designed items are treated equally by the platform in terms of search and recommendation traffic. There is no preferential treatment towards either product type.
[…]
We attribute these improvements to two potential factors. First, the AI-generated digital models enhance the aesthetic appeal of the product images, which can be a challenge for smaller fashion shops. Second, the rapid iteration capabilities of our AI-driven design process significantly reduce the time needed to launch new products from month-level cycles to day-level cycles. This acceleration eliminates the need for human design, prototype production, and model photography, allowing items to remain more closely aligned with fast-changing fashion trends and user preferences.
@@ 6.4. Case Study
This case study demonstrates how PerFusion captures personalized preferences across different merchants. The construction of the text prompt in our application is simply concatenating different descriptive phrases with commas. As shown in Figure 7, the designers from two different e-commerce stores submit the same text prompt for fashion product generation, i.e., “dress,v-neck,slim-fit,long sleeves,long-hair model”. Note that during the training of our model, the descriptive phrases can be shuffled for data augmentation, and the prompt is thereby order-insensitive to phrases. From Figure 7, we draw the following observations:
@@
  • More importantly, even when the text prompt is the same for designers from two different e-commerce stores, the resulting images are tailored to reflect the distinct preferences and records of individual users, with a significant difference in general style and design details. We give our analysis as follows:
@@
This demonstrates that our model can not only satisfy the general aesthetic preference and text-image consistency, but also capture the nuanced personalized preferences across different individuals.
@@ 7. Conclusion
This paper presents an innovative system deployed at Alibaba that leverages personalized text-to-image generation for AI-generated items (AIGI), enabling the “sell it before you make it” business mode. We propose a novel PerFusion framework to model the users’ group-level personalized preferences under such scenarios. We conduct both offline and online experiments to validate the effectiveness of our algorithm. The AI-generated items have achieved over 13% improvement in both click-through rate and conversion rate, as well as 7.9% decrease in return rate, over human-designed ones, highlighting the revolutionary potential of our developed AIGI for e-commerce platforms. Future work could explore more comprehensive personalization from merchants to end customers.
Acknowledgements.
The Shanghai Jiao Tong University team is supported by National Natural Science Foundation of China (624B2096, 62322603). This work was supported by Alibaba Group through Alibaba Research Intern Program.
@@ A.1. Ideal Preference Distribution
The classical pairwise preference optimization algorithms such as DPO (rafailov2024direct; wallace2024diffusion) are derived based on the Bradley-Terry model. In order to consider the preference optmization towards multiple candidates, we have to replace the Bradley-Terry model (bradley1952rank) with the Plackett-Luce model (luce1959individual; plackett1975analysis). Specifically, the Plackett-Luce model probability for a ranking is given by:
@@
$=$ $\sum_{x_{p}\in\mathcal{P}}p(x_{p}\succ x_{n},\forall x_{n}\in\mathcal{N})\;\;\;\;\;\;\;\;\text{(by independent assumption)}$
@@
where $\text{Perm}(\cdot)$ is full permutations of a given set, and, according to previous works (wallace2024diffusion; rafailov2024direct), the reward $r(x,c)$ can be derived as:
@@ A.2. Loss Function
According to previous works on preference optimization (wallace2024diffusion; rafailov2024direct; bai2022training), the training of the target model $p_{\theta}(\cdot)$ is to maximize the log likelihood of the ideal ranking for a given set of candidates. We can first derive the log likelihood of the ideal ranking based on the Plackett-Luce model:
@@
Note that it is intractable to calculate the likelihood of an image, i.e., $p_{\theta}(x|c)$. Therefore, following previous work (wallace2024diffusion), we approximate the denoising trajectory $p_{\theta}(x|c)$ with the diffusion process $q(x|c)$ and derive the final objective: